Please answer and explain thoroughly3. Suppose we were able to measure the amount of oxygen gas formed in units of moles/L, and the rate of formation of oxygen was found to be 0.0125 M/s a. Using the rate law for this reaction and the units associated with each variable, show what the derived units for the rate law constant would be (b) What would be the rate of decomposition of the hydrogen peroxide? Explain your answer.
Answer #1

The rate of decomposition of H2O2 is 1st order

H2O2 ---------->H2O +1/2O2

hence the units for rate constant are /Sec

rate of decomposition of H2O2 is -d(H2O2/dt= d(O2)/2dt

d(O2)/dt= 0.125

-d(H2O2)/dt= 0.125*2= 0.25 M/s